Consider the balanced equation:

2A + 5B ↔ 3C + D

2.0 mol A and 4.0 mol B react to form 1.8 mol C. What is the percent yield of this reaction?
Answer #1

Given reaction:

2A + 4B -----> 3C + D

When 2 moles of A reacts with 5 moles of B it gives 3 moles of C.

Now, if we have 2 moles of C then we need 5 moles of B, however, we have 4 moles of B.

Then the limiting reagent must be B.

for 4 moles of B we need x moles of A.

x = (4*2)/5 = 1.6 moles

Thus A is in excess and B is limiting reagent.

We know 5 moles of B gives 3 moles of C,

then 4 moles of B should give y moles of C,

So , y = (4*3)/5 = 2.4 moles

Thus theoretical yield of C = 2.4 moles = 100%

However, the practical yield = 1.8 moles,

% yield = (1.8*100) / 2.4 = 75%

Thus the percent yield = 75%.